solve this.
3 sin 2x = 2 sin x
1
Expert's answer
2013-03-01T03:50:49-0500
3 sin 2x = 2 sin xLet's expand sin2x:

3*2sin(x)cos(x)-2sin(x)=0Now we factor out common multiplier:

2sin(x)* ( 3cos(x)-1)=0
so sin(x)=0 or cos(x)=1/3Thus, the solutions in general form are

x=pi*k where k integer
or x=2*pi*n+cos^-1(1/3)
or x=2*pi*n-cos^-1(1/3), where n integer